A train travels 150 km in the same time that a plane covers 375 km. If the speed of the plane is 40km per hr less than 3 times the speed of the train, find both speeds.
Answer by stanbon(75887)

A train travels 150 km in the same time that a plane covers 375 km. If the speed of the plane is 40km per hr less than 3 times the speed of the train, find both speeds.
--------------------
Train DATA:
distance = 150 km ; rate = x km/h ; time = 150/x hrs
------------------------
Plane DATA:
distance = 375 km ; rate = 3x-40 km/h ; time = 375/(3x-40) hrs
------------------------
Equation:
time = time
150/x = 375/(3x-40)
150(3x-40) = 375x
450x - 6000 = 375x
75x = 6000
x = 80 km/h (Train rate)
---
3x-40 = 200 km/h (Plane rate)
==================================
Cheers,
Stan H.
=============